OpenStudy (anonymous):

Solving polynomial functions? solve 5x^3 - 6x^2 - 4x - 8 = 0 I kinda understand polynomials, not very much, but I'm extremely confused as how to solve them. If anyone could walk me through this, that'd be great.

OpenStudy (ranga):

Have you been taught the rational roots theorem yet?

OpenStudy (anonymous):

Not that I can recall

OpenStudy (ranga):

Okay. This is a cubic equation and therefore it should have 3 roots (or 3 zeros). Try a few small values for x such as 1, -1, 2, -2, 3, -3 etc. and see if f(x) becomes zero. If it does you would have found one root. Because what is a root. It is the value of x that makes f(x) = 0.

OpenStudy (ranga):

I will do x = 1 and you can try the rest until you find one root. 5x^3 - 6x^2 - 4x - 8 Try x = 1 5(1)^3 - 6(1)^2 - 4(1) - 8 5 - 6 - 4 - 8 = -13 That is not 0 and therefore x = 1 is not a root. Next try x = -1, then x = 2, etc. and let me know as soon as you find the first root.

OpenStudy (anonymous):

2 is a root, because 5(2)^3 -6(2)^2 - 4(2) - 8 40 - 24 - 8 - 8 = 0

OpenStudy (raden):

look at the value of constant, is -8. the factors of -8 is {+-1, +-2, +-4, +-8}. one of or more the values in parantheses can be a factor or a solution for f(x)

OpenStudy (ranga):

Yes! 2 is a root! Do you know synthetic division?

OpenStudy (anonymous):

I've been told to use it but never taught how.

OpenStudy (ranga):

You just found that x = 2 is a root. That means (x-2) is a factor So the original polynomial equation 5x^3 - 6x^2 - 4x - 8 = 0 can be written as: (x-2) * (some other polynomial) = 0 because when you put x = 2, (x-2) becomes 0 and its product will be a 0. 5x^3 - 6x^2 - 4x - 8 = (x-2) * (some other polynomial) To find the "some other polynomial" we need to divide both sides by (x -2) What is (5x^3 - 6x^2 - 4x - 8) / (x-2) ? We have to do either long division or synthetic division to find the quotient. I will do synthetic division here:

OpenStudy (anonymous):

I'm watching some videos about synthetic division and I kind of see how it works. I will try it out myself if you want to save yourself some typing. Thank you for explaining the equation I'd use for synthetic division!

OpenStudy (ranga):

| 2 | 5 -6 -4 -8 | 10 8 8 |___________________________________ 5 4 4 0

OpenStudy (ranga):

The quadratic expression will be: 5x^2 + 4x + 4 (5x^3 - 6x^2 - 4x - 8) = (x - 2)(5x^2 + 4x + 4) See if you can factor 5x^2 + 4x + 4. If not use the quadratic formula and find the two roots.

OpenStudy (ranga):

you are welcome. BTW, you will get two complex roots from the quadratic. And one real root 2 that you found earlier for a total of 3 roots for the cubic equation.

OpenStudy (anonymous):

-2/5 + 4/5i, -2/5 - 4/5i, 2

OpenStudy (ranga):

Yay! You got all three roots correct.